change each equation to slope intercept form and state the slope and y intercept: 6x-4x+8=0
Answer by nerdybill(7384)

Looks like you had a "typo" while you keyed in your problem.
6x-4x+8=0
Perhaps you meant:
6x-4y+8=0 (I'll assume this. If not, please write back)
.
Slope-intercept form is:
y = mx + b
where
m is slope
b is the y-intercept
.
Starting with your equation:
6x-4y+8=0
adding 4y to both sides:
6x+8 = 4y
dividing both sides by 4:
(6/4)x + 2 = y
(3/2)x + 2 = y
flipping it around, we get:
y = (3/2)x + 2
.
slope is 3/2
y-intercept is at (0,2)